If a coin is tossed three times, and then a die is rolled three times, finally a 3-card hand is drawn from a deck of cards, how many different outcomes are possible?
Answer by stanbon(75887)

If a coin is tossed three times: 2^3 = 8 possible outcomes
-----------
and then a die is rolled three times: 6^3 = 216 possible outcomes
-----------
finally a 3-card hand is drawn from a deck of cards: 52C3 = 22100 possible outcomes
=========================
How many different outcomes are possible?
Ans: 8*216*22100 = 38,188,800 different outcomes
======================================================
Cheers,
Stan H.